OpenStudy (anonymous):

What is the simplified form of 4/y^2 - 5/(3y)

OpenStudy (anonymous):

\[\frac{4}{y^2}-\frac{5}{3y}\]?

OpenStudy (anonymous):

yes that is the question these are the options:

OpenStudy (anonymous):

first make the denominators common \[\frac{12}{3y^2}-\frac{5y}{3y^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{12}{3y^2}-\frac{5y}{3y^2}\]solve this :x

OpenStudy (anonymous):

7/ 3y^2

OpenStudy (anonymous):

no 12-5y isn't = 7

OpenStudy (anonymous):

\[\frac{12-5y}{3y^2}\]